OpenStudy (anonymous):

find the sum 1/ b^2c + b/ c^2

OpenStudy (anonymous):

\[\frac{1}{b^2c}+\frac{b}{c^2}\] common denominator will be \(b^2c^2\)

OpenStudy (anonymous):

\[\frac{ 1 }{ b ^{2}c }+\frac{ b }{ c ^{2} }\]

OpenStudy (anonymous):

\[\frac{c}{b^2c^2}+\frac{b^3}{b^2c^2}=\frac{c+b^3}{b^2c^2}\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

please

OpenStudy (anonymous):

\[\frac{ 1 }{ b ^{2}c }+\frac{ b }{ c ^{2} }\] is the original question in order to add fractions, the denominators have to be the same

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

if you wanted to add \[\frac{1}{3^2\times 5}+\frac{3}{5^2}\] what denominator would you have to use?

OpenStudy (anonymous):

\[3^{2}*5^{2}\]

OpenStudy (anonymous):

you need each factor you see to the highest power you see it so the common denominator for \(3^25\) and \(5^2\) would be \[3^2\times 5^2\]

OpenStudy (anonymous):

exactly !

OpenStudy (anonymous):

and you would have to multiply \(\frac{1}{3^2\times 5}\) by \(\frac{5}{5}\) to get \[\frac{5}{3^2\times 5^2}\]

OpenStudy (anonymous):

i have alot of these to do and i have no clue how to do them so do you mind helping me with them and thats where you lost me

OpenStudy (anonymous):

at the last step? lets make sure that is clear

OpenStudy (anonymous):

\[\frac{1}{3^2\times 5}+\frac{3}{5^2}\] we have to write each fraction with the same denominator \(3^2\times 5^2\) right?

OpenStudy (anonymous):

\[\frac{1}{3^2\times 5}\times \frac{5}{5}=\frac{5}{3^2\times 5^2}\\ \frac{3}{5^2}\times \frac{3^2}{3^2}=\frac{3^3}{3^2\times 5^2}\]now that the denominators are the same you can add them

OpenStudy (anonymous):

still lost or did you get that part?

OpenStudy (anonymous):

i will help you with the others, but lets finish this one first

OpenStudy (anonymous):

okay thanks and why do you multiply 5 and 3

OpenStudy (anonymous):

in my example \[\frac{1}{3^2\times 5}\] the denominator is missing a \(5\) to take is \(3^2\times 5^2\) so you have to multiply top and bottom by \(5\)

OpenStudy (anonymous):

ooohh okay that makes sense

OpenStudy (anonymous):

similarly \[\frac{3}{5^2}\] the denominator is missing \(3^2\) so you have to multiply top and bottom by \(3^2\) to build up that fraction

OpenStudy (anonymous):

now we do exactly the same thing, only with \(b\) and \(c\) instead of \(3\) and \(5\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ 1 }{ b ^{2}c }+\frac{ b }{ c ^{2} }\] common denominator is \(b^2\times c^2\)

OpenStudy (anonymous):

the first one multiply top and bottom by \(c\) to get the common denominator

OpenStudy (anonymous):

the second one multiply top and bottom by \(b^2\) to get the common denominator lets cut to the chase

OpenStudy (anonymous):

\[\frac{ 1 }{ b ^{2}c }+\frac{ b }{ c ^{2} }\] \[\frac{c}{b^2c^2}+\frac{b^3}{b^2c^2}=\frac{c+b^3}{b^2c^2}\]

OpenStudy (anonymous):

makes a little more sense

OpenStudy (anonymous):

first fraction multiplied top and bottom by \(c\) second one top and bottom by \(b^2\)

OpenStudy (anonymous):

wanna try another?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

name it

OpenStudy (anonymous):

one sec gotta load the page

OpenStudy (anonymous):

something is wrong with my internet so i will have to get it to work later i will send you a message saying im on

OpenStudy (anonymous):

k good luck